OpenStudy (anonymous):

suppose f(x) = -3x + 2 and g(x) = 1/2x - 1. Find each value. (g(-2))/(f(3))

OpenStudy (anonymous):

g(-2)= 1/2(-2)-1=-1/5 f(3)= -3(3)+2=-7 Therefore, -1/5/-7=1/35

OpenStudy (anonymous):

yay thank you!!!

OpenStudy (anonymous):

Sure